# Beyond Score Changes: Adversarial Attack on No-Reference Image Quality Assessment from Two Perspectives

### v1

@@ Abstract.
No-reference image quality assessment, adversarial attack, multi-objective optimization, correlation-based metrics
copyright: none
@@ 1. Introduction
Deep Neural Networks (DNNs) have demonstrated impressive capabilities in diverse domains, as exemplified by their outstanding performance in a range of studies (Yang et al., 2022; Chen et al., 2023; Gao and Wang, 2023). Among these, Image Quality Assessment (IQA) stands out as a noteworthy application (Li et al., 2020; Yang et al., 2022; Ma et al., 2023; Zhang et al., 2023b). The primary objective of IQA is to predict quality scores for input images, aligning with Mean Opinion Scores (MOS) derived from subjective studies, which provide facilities for the evaluation of multimedia tasks (Li et al., 2021; Zhou and Wang, 2022). IQA models can be categorized into Full-Reference (FR) and No-Reference (NR) types based on the availability of reference images: FR-IQA models have access to the reference image, while NR-IQA models do not (Bosse et al., 2018; Ding et al., 2022). NR-IQA is crucial for designing and optimizing real-world image processing algorithms where the reference image is unavailable, and its growing popularity is attributed to its broader applicability compared to FR-IQA. The evaluation of NR-IQA models is conducted using two major categories of metrics. The first category comprises error-based metrics that measure the prediction error for a single image, such as Mean Squared Error (MSE) and Mean Absolute Error (MAE). The second category is correlation-based, it calculates the correlation relationship between the predicted scores and MOS of a set of images, as observed in metrics like Spearman’s Rank Order Correlation Coefficient (SROCC) (Spearman, 1961), Kendall Rank Order Correlation Coefficient (KROCC) (Kendall, 1938), and Pearson Linear Correlation Coefficient (PLCC) (Pearson, 1895).
Despite these advances, recent research has unveiled a critical issue – the vulnerability of NR-IQA models to adversarial attacks (Zhang et al., 2022; Meftah et al., 2023; Yang et al., 2024; Antsiferova et al., 2024). These attacks added subtle perturbations to input images, thus leading to errors in predicted scores while maintaining the appearance of the perturbed image indistinguishable from the original. We refer to such attacks as error-based attacks. For instance, Zhang et al(2023a) perturbed low/high-quality images to yield high/low predicted scores, which resulted in a significant Root Mean Squared Error (RMSE) between predicted scores before and after the attacks. Besides, some research has concentrated on the score changes across a set of images. Shumitskaya et al(2022) devised a universal perturbation trained for a given set of images to increase their predicted scores.
@@
While error-based attacks effectively underscore the vulnerability of NR-IQA models in the variation of the predicted score, these attacks overlook the impact on correlation-based metrics. As Fig. 1 (left) shows, error-based attacks may randomly increase 10 to 20 points (total 100) to the predicted score of each sample, resulting in a substantial change in predicted scores, but do not significantly alter the relative ranking of these scores. Consequently, such attacks might be ineffective against correlation-based metrics like SROCC. Merely targeting the error-based metrics of NR-IQA models is insufficient for a comprehensive attack strategy. An effective attack should take into account all evaluation metrics of the model, including both error-based metrics and correlation-based metrics. Furthermore, it is important to note that the NR-IQA task is a regression problem where the correlation, like ranking correlation, plays a significant role in its applications. This is particularly evident in applications such as image retrieval (Somasundaran et al., 2020; Li et al., 2022). Perturbations that affect ranking could seriously compromise retrieval accuracy, underscoring the need to analyze model robustness from the viewpoint of ranking correlation. In light of these insights, considering attacks that disrupt both error-based and correlation-based metrics becomes a pressing need.
@@
After the problem formulation, We consider how to solve this problem. Directly optimizing MSE and SROCC on the whole image set is challenging due to its high dimensionality. Instead of it, we propose a two-stage SROCC-MSE Attack (SMA) to generate adversarial examples. In Stage One, our objective is to identify optimal target scores that significantly reduce SROCC and increase MSE between the predicted scores of attacked and clean images. As SROCC is non-differentiable due to the discrete rank function, we approximate it using a differential function and employ the Lagrangian multiplier method to solve the optimization problem. In Stage Two, we utilize the fast gradient sign method (Goodfellow et al., 2015) to generate adversarial examples, striving to make their predicted scores as close as possible to the target scores identified in Stage One. While exact equivalence between the predicted and target scores may not be achievable, our experiments provide compelling evidence that the discrepancy between them is small.
We conduct experiments on the CLIVE dataset (Ghadiyaram and Bovik, 2016), which includes authentic distorted images sourced from the wild and is challenging for NR-IQA. Attacks are conducted on four widely-used NR-IQA models, DBCNN (Zhang et al., 2020), HyperIQA (Su et al., 2020), MANIQA (Yang et al., 2022), and LIQE (Zhang et al., 2023b). Our approach is compared with four existing attack methods, P-attack (Zhang et al., 2022), OUAP (Shumitskaya et al., 2024), K-attack (Korhonen and You, 2022), and UAP (Shumitskaya et al., 2022). Our evaluations spanned seven metrics, considering both error and correlation perspectives. The results demonstrate that SMA offers a powerful attack that influences not only the prediction score of individual images but also the correlation within the set of predicted scores. For instance, in all attacked models, SMA significantly disrupts the SROCC to negative values while simultaneously maintaining a substantial change in the scores of individual images. It demonstrates the effectiveness of SMA within our correlation-error-based attacks framework. The findings underscore the vulnerability of NR-IQA in maintaining individual scores and correlations, paving the way for further research on developing more secure and robust NR-IQA models.
@@ 2.1. Attacks on NR-IQA Models
Perceptual attack (Zhang et al., 2022) is proposed to attack available NR-IQA models. It changes predicted scores to a certain extent while maintaining the similarity between clean images and adversarial examples, incorporating human feedback in the loop. Korhonen and You (2022) developed a neural network capable of generating adversarial examples, which can be used to attack unknown models. In contrast to the perceptual attack, the universal adversarial perturbation (UAP) (Shumitskaya et al., 2022) is designed to execute attacks by considering a set of images. UAP aimed at generating a universal perturbation applicable to a set of images, to collectively increase the predicted scores by an NR-IQA model. It implies the adversarial perturbation remains the same for all input images within the set. Similar to UAP, OUAP (2024) also generates a universal perturbation for a set of images, maximizing the score predicted by the attacked NR-IQA model on a set of images. However, the imperceptibility of universal perturbations is not satisfactory.
@@ 2.2. Differential Ranking
(1) $r:\boldsymbol{s}\rightarrow\text{Perm}(n),$
@@
Many studies have explored alternative approaches to handle differentiable proxies for ranking. One feasible way is directly approximating ranking metrics, such as minimizing a smooth approximation of ranking measures through gradient descent strategies (Chapelle and Wu, 2010) or employing diverse forms of loss-augmented inference (Mohapatra et al., 2018). Another prevalent method is referred to as “soft” ranking, which can be integrated into any differential loss function. Various techniques have been employed to achieve this goal. For instance, Taylor et al(2008) introduced a method where each score is smoothed using equal variance Gaussian distributions. Engilberge et al(2019) took a different approach by training a deep neural network to learn the entire sorting operation. However, these methods tend to be time-consuming. In contrast, Blondel et al(2020) proposed an efficient approach by constructing differentiable operators as projections onto the convex hull of permutations, demonstrating speed and computational friendliness. In this paper, we adopt the method proposed by Blondel et al(2020) during the SROCC optimization phase.
@@ 3.1. Problem Definition
(2) $\begin{split}\min_{\mathcal{I}^{\prime}}~{}~{}\mathcal{L}_{\text{cor}}&(f(\mathcal{I}^{\prime}),f(\mathcal{I}))-\lambda\mathcal{L}_{\text{err}}(f(\mathcal{I}^{\prime}),f(\mathcal{I})),\\ \text{s.t. }&D(I^{\prime}_{j},I_{j})\leqslant\epsilon,\forall j\in\{1,\cdots,N\}.\end{split}$
@@
In this paper, we focus on the $\mathcal{L}_{\text{cor}}(\cdot,\cdot)$ related to the ranking, and mainly consider the SROCC (Spearman, 1961) which is commonly used in IQA tasks (Siniukov et al., 2022; Madhusudana et al., 2022; Siniukov et al., 2023), i.e.,
(3) $\mathcal{L}_{\text{cor}}(\boldsymbol{s}^{\prime},\boldsymbol{s})$ $=\text{SROCC}(\boldsymbol{s}^{\prime},\boldsymbol{s})$
@@
where $r_{j}(\cdot)$ is the $j^{th}$ element of the ranking function $r(\cdot)$. For $\mathcal{L}_{\text{err}}(\cdot,\cdot)$, we leverage MSE to measure the error between $s^{\prime}_{j}$ and $s_{j}$:
(4) $\mathcal{L}_{\text{err}}(\boldsymbol{s}^{\prime},\boldsymbol{s})=\text{MSE}(\boldsymbol{s}^{\prime},\boldsymbol{s})=\frac{1}{N}\sum_{j=1}^{N}\left(s^{\prime}_{j}-s_{j}\right)^{2}.$
The perceptual distance $D(\cdot,\cdot)$ has many choices (Zhang et al., 2022), such as the $\ell_{\infty}$ norm, SSIM (Wang et al., 2004), LPIPS (Zhang et al., 2018), and so on (Sheikh and Bovik, 2006; Reisenhofer et al., 2018; Ding et al., 2022). Among them, we choose the most commonly used $\ell_{\infty}$ norm as $D(\cdot,\cdot)$ for its simplicity in computation and optimization.
@@
(5) $\begin{split}\min_{\mathcal{I}^{\prime}}~{}~{}\text{SROCC}&(f(\mathcal{I}^{\prime}),f(\mathcal{I}))-\lambda_{\text{MSE}}\text{MSE}(f(\mathcal{I}^{\prime}),f(\mathcal{I})),\\ \text{s.t. }&\left\|I^{\prime}_{j}-I_{j}\right\|_{\infty}\leqslant\epsilon,\forall j\in\{1,\cdots,N\},\end{split}$
@@ 3.3. Target Scores Optimization
(6) $\boldsymbol{s}^{*}=\mathop{\arg\min}_{\boldsymbol{\tilde{s}}}\text{SROCC}(\boldsymbol{\tilde{s}},\boldsymbol{s})-\lambda_{\text{MSE}}\text{MSE}(\boldsymbol{\tilde{s}},\boldsymbol{s}),$
@@
(7) $\boldsymbol{s}^{*}=\mathop{\arg\min}_{\boldsymbol{\tilde{s}}}\left(\text{SROCC}(\boldsymbol{\tilde{s}},\boldsymbol{s})-\lambda_{\text{Var}}\text{Var}(\boldsymbol{\tilde{s}})-\lambda_{\text{MSE}}\text{MSE}(\boldsymbol{\tilde{s}},\boldsymbol{s})\right),$
@@
(8) $\text{Var}(\boldsymbol{\tilde{s}})=\frac{1}{N-1}\sum_{j=1}^{N}\left(\tilde{s}_{j}-\frac{1}{N}\sum_{k=1}^{N}\tilde{s}_{k}\right)^{2}$
@@
To solve the optimization problem defined in Eq. (7) with gradient-based methods, we approximate the non-differential ranking function $r(\cdot)$ with a differential function $r^{\beta}(\cdot)$ where $\beta>0$ serves as a hyper-parameter. Following the approach of Blondel et al(2020), $r(\cdot)$ is approximated by
(9) $r^{\beta}(\boldsymbol{s})=\mathop{\arg\min}_{\boldsymbol{z}\in\textbf{conv}[N]}\frac{1}{2}\left\|\boldsymbol{z}+\frac{\boldsymbol{s}}{\beta}\right\|^{2},$
where $\textbf{conv}[N]$ is a convex polytope whose vertices correspond to permutations in $[N]$. The notation $[N]$ represents all permutations of $\{1,\dots,N\}$ which contain $N!$ vertices111[N]={(1,2,3…,N),(2,1,3,…,N),…,(N,N-1,N-2…,1)}.. As illustrated in Fig. 3, $r^{\beta}(\boldsymbol{s})$ calculates the Euclidean projection of $\boldsymbol{s}/\beta$ on to $\textbf{conv}[N]$. If the projection of $\boldsymbol{s}/\beta$ lies outside $\textbf{conv}[N]$, $r^{\beta}(\boldsymbol{s})$ computes the nearest vertex of $\textbf{conv}[N]$ to the projection. The function $r^{\beta}(\cdot)$ is differential, as proven in the work (Blondel et al., 2020). Furthermore, the approximation error, measured by $\|r^{\beta}(\boldsymbol{s})-r(\boldsymbol{s})\|_{2}^{2}$, is close to zero when $\beta$ is sufficiently small (Blondel et al., 2020). As a special case in (2020), we have:
@@
(10) $\beta\leqslant\min_{i\in\{1,\dots,N-1\}}\boldsymbol{s}_{r^{-1}(i)}-\boldsymbol{s}_{r^{-1}(i+1)}.$
@@
(11) $\min_{\tilde{s}}1-\frac{6}{N(N^{2}-1)}\sum_{j=1}^{N}\left(r^{\beta}_{j}(\boldsymbol{\tilde{s}})-r^{\beta}_{j}(\boldsymbol{s})\right)^{2}$

### v2

@@ 1. Introduction
Deep Neural Networks (DNNs) have demonstrated impressive capabilities in diverse domains, as exemplified by their outstanding performance in a range of studies Yang et al. (2022); Chen et al. (2023); Gao and Wang (2023). Among these, Image Quality Assessment (IQA) stands out as a noteworthy application Li et al. (2020); Yang et al. (2022); Ma et al. (2023); Zhang et al. (2023b). The primary objective of IQA is to predict quality scores for input images, aligning with Mean Opinion Scores (MOS) derived from subjective studies, which provide facilities for the evaluation of multimedia tasks Li et al. (2021); Zhou and Wang (2022). IQA models can be categorized into Full-Reference (FR) and No-Reference (NR) types based on the availability of reference images: FR-IQA models have access to the reference image, while NR-IQA models do not Bosse et al. (2018); Ding et al. (2022). NR-IQA is crucial for designing and optimizing real-world image processing algorithms where the reference image is unavailable, and its growing popularity is attributed to its broader applicability compared to FR-IQA. The evaluation of NR-IQA models is conducted using two major categories of metrics. The first category comprises error-based metrics that measure the prediction error for a single image, such as Mean Squared Error (MSE) and Mean Absolute Error (MAE). The second category is correlation-based, it calculates the correlation relationship between the predicted scores and MOS of a set of images, as observed in metrics like Spearman’s Rank Order Correlation Coefficient (SROCC) Spearman (1961), Kendall Rank Order Correlation Coefficient (KROCC) Kendall (1938), and Pearson Linear Correlation Coefficient (PLCC) Pearson (1895).
Despite these advances, recent research has unveiled a critical issue – the vulnerability of NR-IQA models to adversarial attacks Zhang et al. (2022); Meftah et al. (2023); Yang et al. (2024); Antsiferova et al. (2024). These attacks added subtle perturbations to input images, thus leading to errors in predicted scores while maintaining the appearance of the perturbed image indistinguishable from the original. We refer to such attacks as error-based attacks. For instance, Zhang et alZhang et al. (2023a) perturbed low/high-quality images to yield high/low predicted scores, which resulted in a significant Root Mean Squared Error (RMSE) between predicted scores before and after the attacks. Besides, some research has concentrated on the score changes across a set of images. Shumitskaya et alShumitskaya et al. (2022) devised a universal perturbation trained for a given set of images to increase their predicted scores.
@@
While error-based attacks effectively underscore the vulnerability of NR-IQA models in the variation of the predicted score, these attacks overlook the impact on correlation-based metrics. As Fig. 1 (left) shows, error-based attacks may randomly increase 10 to 20 points (total 100) to the predicted score of each sample, resulting in a substantial change in predicted scores, but do not significantly alter the relative ranking of these scores. Consequently, such attacks might be ineffective against correlation-based metrics like SROCC. Merely targeting the error-based metrics of NR-IQA models is insufficient for a comprehensive attack strategy. An effective attack should take into account all evaluation metrics of the model, including both error-based metrics and correlation-based metrics. Furthermore, it is important to note that the NR-IQA task is a regression problem where the correlation, like ranking correlation, plays a significant role in its applications. This is particularly evident in applications such as image retrieval Somasundaran et al. (2020); Li et al. (2022). Perturbations that affect ranking could seriously compromise retrieval accuracy, underscoring the need to analyze model robustness from the viewpoint of ranking correlation. In light of these insights, considering attacks that disrupt both error-based and correlation-based metrics becomes a pressing need.
@@
After the problem formulation, We consider how to solve this problem. Directly optimizing MSE and SROCC on the whole image set is challenging due to its high dimensionality. Instead of it, we propose a two-stage SROCC-MSE Attack (SMA) to generate adversarial examples. In Stage One, our objective is to identify optimal target scores that significantly reduce SROCC and increase MSE between the predicted scores of attacked and clean images. As SROCC is non-differentiable due to the discrete rank function, we approximate it using a differential function and employ the Lagrangian multiplier method to solve the optimization problem. In Stage Two, we utilize the fast gradient sign method Goodfellow et al. (2015) to generate adversarial examples, striving to make their predicted scores as close as possible to the target scores identified in Stage One. While exact equivalence between the predicted and target scores may not be achievable, our experiments provide compelling evidence that the discrepancy between them is small.
We conduct experiments on the CLIVE dataset Ghadiyaram and Bovik (2016), which includes authentic distorted images sourced from the wild and is challenging for NR-IQA. Attacks are conducted on four widely-used NR-IQA models, DBCNN Zhang et al. (2020), HyperIQA Su et al. (2020), MANIQA Yang et al. (2022), and LIQE Zhang et al. (2023b). Our approach is compared with four existing attack methods, P-attack Zhang et al. (2022), OUAP Shumitskaya et al. (2024), K-attack Korhonen and You (2022), and UAP Shumitskaya et al. (2022). Our evaluations spanned seven metrics, considering both error and correlation perspectives. The results demonstrate that SMA offers a powerful attack that influences not only the prediction score of individual images but also the correlation within the set of predicted scores. For instance, in all attacked models, SMA significantly disrupts the SROCC to negative values while simultaneously maintaining a substantial change in the scores of individual images. It demonstrates the effectiveness of SMA within our correlation-error-based attacks framework. The findings underscore the vulnerability of NR-IQA in maintaining individual scores and correlations, paving the way for further research on developing more secure and robust NR-IQA models.
@@ 2.1. Attacks on NR-IQA Models
Perceptual attack Zhang et al. (2022) is proposed to attack available NR-IQA models. It changes predicted scores to a certain extent while maintaining the similarity between clean images and adversarial examples, incorporating human feedback in the loop. Korhonen and You Korhonen and You (2022) developed a neural network capable of generating adversarial examples, which can be used to attack unknown models. In contrast to the perceptual attack, the universal adversarial perturbation (UAP) Shumitskaya et al. (2022) is designed to execute attacks by considering a set of images. UAP aimed at generating a universal perturbation applicable to a set of images, to collectively increase the predicted scores by an NR-IQA model. It implies the adversarial perturbation remains the same for all input images within the set. Similar to UAP, OUAP Shumitskaya et al. (2024) also generates a universal perturbation for a set of images, maximizing the score predicted by the attacked NR-IQA model on a set of images. However, the imperceptibility of universal perturbations is not satisfactory.
@@ 2.2. Differential Ranking
$r:\boldsymbol{s}\rightarrow\text{Perm}(n),$ (1)
@@
Many studies have explored alternative approaches to handle differentiable proxies for ranking. One feasible way is directly approximating ranking metrics, such as minimizing a smooth approximation of ranking measures through gradient descent strategies Chapelle and Wu (2010) or employing diverse forms of loss-augmented inference Mohapatra et al. (2018). Another prevalent method is referred to as “soft” ranking, which can be integrated into any differential loss function. Various techniques have been employed to achieve this goal. For instance, Taylor et alTaylor et al. (2008) introduced a method where each score is smoothed using equal variance Gaussian distributions. Engilberge et alEngilberge et al. (2019) took a different approach by training a deep neural network to learn the entire sorting operation. However, these methods tend to be time-consuming. In contrast, Blondel et alBlondel et al. (2020) proposed an efficient approach by constructing differentiable operators as projections onto the convex hull of permutations, demonstrating speed and computational friendliness. In this paper, we adopt the method proposed by Blondel et alBlondel et al. (2020) during the SROCC optimization phase.
@@ 3.1. Problem Definition
$\begin{split}\min_{\mathcal{I}^{\prime}}\leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{L}_{\text{cor}}&(f(\mathcal{I}^{\prime}),f(\mathcal{I}))-\lambda\mathcal{L}_{\text{err}}(f(\mathcal{I}^{\prime}),f(\mathcal{I})),\\ \text{s.t. }&D(I^{\prime}_{j},I_{j})\leqslant\epsilon,\forall j\in\{1,\cdots,N\}.\end{split}$ (2)
@@
In this paper, we focus on the $\mathcal{L}_{\text{cor}}(\cdot,\cdot)$ related to the ranking, and mainly consider the SROCC Spearman (1961) commonly used in IQA tasks Siniukov et al. (2022); Madhusudana et al. (2022); Siniukov et al. (2023), i.e.,
$\mathcal{L}_{\text{cor}}(\boldsymbol{s}^{\prime},\boldsymbol{s})$ $=\text{SROCC}(\boldsymbol{s}^{\prime},\boldsymbol{s})$ (3)
@@
Here, $r_{j}(\cdot)$ is the $j^{th}$ element of the ranking function $r(\cdot)$. For $\mathcal{L}_{\text{err}}(\cdot,\cdot)$, we leverage MSE to measure the error between $s^{\prime}_{j}$ and $s_{j}$:
$\mathcal{L}_{\text{err}}(\boldsymbol{s}^{\prime},\boldsymbol{s})=\text{MSE}(\boldsymbol{s}^{\prime},\boldsymbol{s})=\frac{1}{N}\sum_{j=1}^{N}\left(s^{\prime}_{j}-s_{j}\right)^{2}.$ (4)
The perceptual distance $D(\cdot,\cdot)$ has many choices Zhang et al. (2022), such as the $\ell_{\infty}$ norm, SSIM Wang et al. (2004), LPIPS Zhang et al. (2018), and so on Sheikh and Bovik (2006); Reisenhofer et al. (2018); Ding et al. (2022). Among them, we choose the most commonly used $\ell_{\infty}$ norm as $D(\cdot,\cdot)$ for its simplicity in computation and optimization.
@@
$\begin{split}\min_{\mathcal{I}^{\prime}}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{SROCC}&(f(\mathcal{I}^{\prime}),f(\mathcal{I}))-\lambda_{\text{MSE}}\text{MSE}(f(\mathcal{I}^{\prime}),f(\mathcal{I})),\\ \text{s.t. }&\left\|I^{\prime}_{j}-I_{j}\right\|_{\infty}\leqslant\epsilon,\forall j\in\{1,\cdots,N\},\end{split}$ (5)
@@ 3.3. Target Scores Optimization
$\boldsymbol{s}^{*}=\mathop{\arg\min}_{\boldsymbol{\tilde{s}}}\text{SROCC}(\boldsymbol{\tilde{s}},\boldsymbol{s})-\lambda_{\text{MSE}}\text{MSE}(\boldsymbol{\tilde{s}},\boldsymbol{s}),$ (6)
@@
$\boldsymbol{s}^{*}=\mathop{\arg\min}_{\boldsymbol{\tilde{s}}}\left(\text{SROCC}(\boldsymbol{\tilde{s}},\boldsymbol{s})-\lambda_{\text{Var}}\text{Var}(\boldsymbol{\tilde{s}})-\lambda_{\text{MSE}}\text{MSE}(\boldsymbol{\tilde{s}},\boldsymbol{s})\right),$ (7)
@@
$\text{Var}(\boldsymbol{\tilde{s}})=\frac{1}{N-1}\sum_{j=1}^{N}\left(\tilde{s}_{j}-\frac{1}{N}\sum_{k=1}^{N}\tilde{s}_{k}\right)^{2}$ (8)
@@
To solve the optimization problem defined in Eq. (7) with gradient-based methods, we approximate the non-differential ranking function $r(\cdot)$ with a differential function $r^{\beta}(\cdot)$ where $\beta>0$ serves as a hyper-parameter. Following the approach of Blondel et alBlondel et al. (2020), $r(\cdot)$ is approximated by
$r^{\beta}(\boldsymbol{s})=\mathop{\arg\min}_{\boldsymbol{z}\in\textbf{conv}[N]}\frac{1}{2}\left\|\boldsymbol{z}+\frac{\boldsymbol{s}}{\beta}\right\|^{2},$ (9)
where $\textbf{conv}[N]$ is a convex polytope whose vertices correspond to permutations in $[N]$. The notation $[N]$ represents all permutations of $\{1,\dots,N\}$ which contain $N!$ vertices111[N]={(1,2,3…,N),(2,1,3,…,N),…,(N,N-1,N-2…,1)}.. As illustrated in Fig. 3, $r^{\beta}(\boldsymbol{s})$ calculates the Euclidean projection of $\boldsymbol{s}/\beta$ on to $\textbf{conv}[N]$. If the projection of $\boldsymbol{s}/\beta$ lies outside $\textbf{conv}[N]$, $r^{\beta}(\boldsymbol{s})$ computes the nearest vertex of $\textbf{conv}[N]$ to the projection. The function $r^{\beta}(\cdot)$ is differential, as proven in the work Blondel et al. (2020). Furthermore, the approximation error, measured by $\|r^{\beta}(\boldsymbol{s})-r(\boldsymbol{s})\|_{2}^{2}$, is close to zero when $\beta$ is sufficiently small Blondel et al. (2020). As a special case in Blondel et al. (2020), we have:
@@
$\beta\leqslant\min_{i\in\{1,\dots,N-1\}}\boldsymbol{s}_{r^{-1}(i)}-\boldsymbol{s}_{r^{-1}(i+1)}.$ (10)
@@
$\min_{\tilde{s}}1-\frac{6}{N(N^{2}-1)}\sum_{j=1}^{N}\left(r^{\beta}_{j}(\boldsymbol{\tilde{s}})-r^{\beta}_{j}(\boldsymbol{s})\right)^{2}$ (11)